how to solve i^15 step by step
Answer by jim_thompson5910(35256)

i^15

i^(12+3)

i^(12)*i^3

(i^(3*4))*i^3

(i^3)^4*i^3

(-i)^4*(-i)

(-1)^4*(i)^4*(-i)

(-1)^4*(i^2)^2*(-i)

(-1)^4*(-1)^2*(-i)

(-1)^4*1*(-i)

(-1)^4*(-i)

1*(-i)

-i

So i^15 = -i

Shortcut: divide 15 by 4 to get 3 remainder 3. The remainder of 3 means that i^(15) = i^3. Since i^3 = -i, this means i^15 = -i